3. An insulated tank is initially divided with a membrane. See Fig. 3. One portion has a volume of 1 m3 and contains air at 11 2 Imal 3m3 6 bar ibar 1500 1500 membrar Fig. 3.
Answer #1

Given GO 1.013 kulty K, CU = 0.726 KS ., kan K= 1.345 9.k 0 كيم كا N2 = 3m3 Vi= 1 m3 Ti = 150°c T2 = 150°c Pi= 6 bar P2= 1 baet portion: By Iebed gog Qualion, Piui= MRT Ra e-Coc 1.0l – 0.42= 0,283 k fait 6x10) 4 = mixo.287 x 423 m= 4.942 kg portion:final pressure Cf) = 2.249 bas. Entropy dange, (AS) suston e 60s)+(45)2 + (As) et (29) = m(Gp dn - R en € 4202 (1.013 enero )